OpenStudy (anonymous):

Write the equation of the line that passes through (–1, 5) with a slope of 2 in standard form. –2x – y = –7 2x – y = –7 2x – y = 11 –2x + y = –11

OpenStudy (anonymous):

Use point-slope form: y - y1 = m(x - x1) y - 5 = 2(x + 1) y - 5 = 2x + 2 Then, rearrange the terms into standard form Ax + By = C 2x - y = -7